OpenStudy (sugarmochi):

Solve the following equation: 3x^2 - 11x - 4 = 0

OpenStudy (naka354):

3*x^2 - 11*x - 4 = 0 (3x + 1)(x - 4) = 0 etc

OpenStudy (sugarmochi):

where would the 11 have gone-- ?

OpenStudy (3mar):

May I help?

OpenStudy (mrnood):

This is called factorisation - the 2 brackets will multip[ly out to be the SAME as your original equation (including the -11x) What methods do you know for solving quadratics? The factorising is the easiest

OpenStudy (sugarmochi):

Uhm- factorising- I know the basics but i dont exactly know how to do it here

OpenStudy (naka354):

General form of quadratic equation: ax^2 + bx + c = 0 <=> (x - p)(x - q) = 0 has roots p and q. (x - p)(x - q) = x² - (p+q)x + p*q = 0 p + q = -b/a p*q = c/a Go on please analogically.

OpenStudy (mrnood):

@naka354 Has given you the factorisation above: (3x + 1)(x - 4) = 0 The 2 answers follow directly from that

OpenStudy (sugarmochi):

I'm still really r e a l l y confused on the process of how 3x^2 - 11x - 4 = 0 became (3x + 1)(x - 4) = 0 --

OpenStudy (mrnood):

If you multiply out the brackets using FOIL then you will see they ar the same